$14, payable debt earnings $131 depreciation )
Answer #1

Calculation of Accumulated Depreciation for 2017= $141,400+ 46,000 = 187,400

Retained Earnings 2016= Total Assets- Total Liabilites- Common Stock

=401200-223,000-131,900

=46,300

Retained Earnings 2017= Total Assets- Total Liabilites- Common Stock

=440,900-198,000-181,800

=61,100